# Optimal Design of Automated Market Markers on Decentralized Exchanges

### v2

@@ 3.3.1 LP’s Investment Return on DEX
$=$ $\underbrace{\frac{p_{k}^{A}y_{(k,1)}^{A}+p_{k}^{B}y_{(k,1)}^{B}}{p_{k}^{A}y_{(k,0)}^{A}+p_{k}^{B}y_{(k,0)}^{B}}}_{R^{M}_{(k+1,1)}}\cdot\underbrace{\frac{p_{k}^{A}y_{(k,2)}^{A}+p_{k}^{B}y_{(k,2)}^{B}}{p_{k}^{A}y_{(k,1)}^{A}+p_{k}^{B}y_{(k,1)}^{B}}}_{R^{M}_{(k+1,2)}}\cdot\underbrace{\frac{p_{k+1}^{A}y_{(k,2)}^{A}+p_{k+1}^{B}y_{(k,2)}^{B}}{p_{k}^{A}y_{(k,2)}^{A}+p_{k}^{B}y_{(k,2)}^{B}}}_{R^{M}_{(k+1,3)}}\cdot\underbrace{\frac{p_{k+1}^{A}y_{(k,3)}^{A}+p_{k+1}^{B}y_{(k,3)}^{B}}{p_{k+1}^{A}y_{(k,2)}^{A}+p_{k+1}^{B}y_{(k,2)}^{B}}}_{R^{M}_{(k+1,4)}},$ (23)
where $R^{M}_{(k+1,1)}$, $R^{M}_{(k+1,2)}$, and $R^{M}_{(k+1,4)}$ stand for the change of the liquidity pool value due to trading by investors and arbitrageurs in Steps 2, 3, and 5, respectively, and $R^{M}_{(k+1,3)}$ stands for the change of the liquidity pool value due to the price shock in Step 4. By (4), (17), (19), and (21), we derive
@@ 3.4.2 Dynamic Programming
Recall that starting from any exchange rate ratio $S_{k}$ at time $k$, $S_{k+1}$ always falls in $[1/(1+f),1+f]$. Thus, we assume without loss of generality that the state variable $S_{k}$ takes values in $[1/(1+f),1+f]$. Denote by
${\cal S}:=[1/(1+f),1+f]\times(0,\infty)$
@@
    $\delta\bar{R}^{1-\gamma}<1,$ (52)
    where
    $\bar{R}:=\begin{cases}\sup_{(s,z)\in{\cal S}}\sup_{\boldsymbol{\omega}_{k}\in\Omega}U^{-1}\left(\mathbb{E}\Big{[}U\big{(}R^{p}_{k+1}\big{)}\big{]}\mid S_{k}=s,Z_{k}=z\Big{]}\right),&\gamma\in(0,1),\\ \inf_{(s,z)\in{\cal S}}\sup_{\boldsymbol{\omega}_{k}\in\Omega}U^{-1}\left(\mathbb{E}\Big{[}U\big{(}R^{p}_{k+1}\big{)}\big{]}\mid S_{k}=s,Z_{k}=z\Big{]}\right),&\gamma>1.\end{cases}$ (53)
@@ 3.4.3 The Case of CGMMM Pricing Function
Combining (66)-(69) with (23) and (62)–(65), we derive the return on AMM: $R^{M}_{k+1}/R^{B}_{k+1}$ is a function of $S_{k}$ and random disturbances $(\xi_{k+1},I_{k+1},R^{A}_{k+1}/R^{B}_{k+1})$, but it is independent of $Z_{k}$.
Therefore, we conclude that for the CGMMM pricing function, given the exchange ratio $S_{k}$, the fundamental exchange rate $Z_{k}$ is irrelevant for the LP’s decision. Thus, we can reduce the dimension of the state space.
It is also worth mentioning that for the CGMMM pricing function, the exchange rate ratio takes the form
$S_{k}=\frac{\eta}{(1-\eta)}\times\frac{y_{k}^{B}p_{k}^{B}}{y_{k}^{A}p_{k}^{A}},$
where $(y_{k}^{B}p_{k}^{B})/(y_{k}^{A}p_{k}^{A})$ is the relative portfolio weight of asset $B$ to asset $A$ in the liquidity pool. In view of Proposition 3.1, the no-trading region for arbitrageurs can then be written as an interval for this relative portfolio weight. This explains the no-fee interval assumed in Evans et al. (2021).
@@ 4.1 Data Description and Parameter Estimation
We set the length of one period to be 8 hours and assume the LP consumes every 24 hours, so that $N=3$ in our setting. We set the LP’s risk aversion $\gamma=2$. We take ETH and BTC as assets $A$ and $B$, respectively, and estimate the market parameters $\mu_{A}$, $\mu_{B}$, $\sigma_{A}$, $\sigma_{B}$, and $\rho$ using the 8-hour trading data of ETH/USDT and BTC/USDT from July 1st, 2022 to June 30, 2023 obtained from Binance Market Data (https://data.binance.vision). We set $\eta=0.5$ corresponding to the CPMM used in Uniswap, $f=0.5\%$ in the middle of the range of the fee levels on Uniswap v3 (0.05%, 0.3%, and 1%), $\sigma_{I}=0.02$ and $\alpha=0.5$. For convenience, we summarize the parameter values in Table 1.
@@
Market Parameters
$\delta$ 0.998 $R_{f}$ 1+0.002%
$\mu_{A}$ 0.05% $\mu_{B}$ 0.038%
$\sigma_{A}$ 1.99% $\sigma_{B}$ 1.52%
$\rho$ 0.8642 $\alpha$ 0.5
$\sigma_{I}$ 2%
LP’s Attributes
$N$ 3 $\gamma$ 2
AMM Parameters
$\eta$ 0.5 $f$ 0.5%
@@ 4.2.2 Price Slippage
The price slippage is the difference between the marginal exchange rate on AMM and the average exchange rate of trading a positive amount of an asset on AMM. Under CGMMM, it is straightforward to derive that the relative price slippage for acquiring $D_{A}>0$ amount of asset $A$ at marginal exchange rate $E_{0}$ is
$\frac{-D^{B}/D^{A}-E_{0}}{E_{0}}=\frac{\frac{1-\eta}{\eta}\left(\big{(}1-D^{A}/y^{A}\big{)}^{\eta/(\eta-1)}-1\right)}{D^{A}/y^{A}},$
where $D^{B}$ is determined by the pricing equation (1), and therefore $-D^{B}/D^{A}$ stands for the average exchange rate. It is straightforward to verify that the above relative price slippage is strictly increasing in the trading size $D^{A}/y^{A}$ and in $\eta$. Similarly, the relative price slippage for acquiring $D_{B}>0$ amount of asset $B$ at marginal exchange rate $E_{0}^{\prime}$ is
$\frac{-D^{A}/D^{B}-E_{0}^{\prime}}{E_{0}^{\prime}}=\frac{\frac{\eta}{1-\eta}\left(\big{(}1-D^{B}/y^{B}\big{)}^{(1-\eta)/\eta}-1\right)}{D^{B}/y^{B}},$
which is strictly increasing in the trading size $D^{B}/y^{B}$ and strictly decreasing in $\eta$.
@@ 4.2.3 LP’s Fee Revenue from Liquidity Traders and Loss to Arbitrageurs
$\approx\left[\eta R^{A}(1+f\mathbf{1}_{\tilde{R}^{-1}>(1+f)})\bar{\varphi}^{A}(\tilde{R}^{-1})+(1-\eta)R^{B}(1+f\mathbf{1}_{\tilde{R}^{-1}<1/(1+f)})\bar{\varphi}^{B}(\tilde{R}^{-1})\right]$
$\quad\cdot(y^{A}p^{A}+y^{B}p^{B}),$
@@
Milionis, Moallemi, Roughgarden and Zhang (2024) consider a continuous-time setting with asset $B$ as the numéraire and thus $R^{B}\equiv 1$. In this setting, $\tilde{R}$ represents the fundamental exchange rate change in an infinitesimally small period of time and thus has a mean of 1, and the expectation of $(\tilde{R}-1)^{2}$ is proportional the fundamental exchange rate volatility. Thus, Milionis, Moallemi, Roughgarden and Zhang (2024) derive that the LP’s expected loss to arbitrageurs in their setting is proportional to $\eta(1-\eta)$ and the volatility of the fundamental exchange rate; see Example 2 therein.
@@
$\approx\left[\eta\big{(}\left(I^{-1}/(1+f)\right)^{1-\eta}-1\big{)}\mathbf{1}_{I^{-1}>1+f}+(1-\eta)\big{(}\left(I^{-1}(1+f)\right)^{-\eta}-1\big{)}\mathbf{1}_{I^{-1}<1/(1+f)}\right]$
$\quad\cdot(p^{A}y^{A}+p^{B}y^{B})f\xi,$
@@ 4.5 Empirical Test
We use data in 110 AMM pools on Uniswap v3333See https://thegraph.com/hosted-service/subgraph/uniswap/uniswap-v3. and asset price data on Binance [^4] See https://data.binance.vision/. from July 1, 2022 to June 30, 2023. For each AMM pool $j$, we collect its unit trading fee $f_{j}$ from Uniswap v3 and estimate the volatilities of the two assets in the pool $\sigma_{A,j}$ and $\sigma_{B,j}$ and their correlation $\rho_{j}$ using the Binance asset price data. Then we estimate the volatility of the fundamental exchange rate for the pool $j$ as
$\sigma_{j}=\sqrt{\sigma_{A,j}^{2}-2\rho_{j}\sigma_{A,j}\sigma_{B,j}+\sigma_{B,j}^{2}}$
We then run a linear regression between the unit trading fee $f$ and the corresponding fundamental exchange rate volatility $\sigma$, asset correlation $\rho$, and asset $B$’s volatility $\sigma_{B}$ across the 110 AMM pools, and the results are reported in Table 2. At the 5% significance level, the unit trading fee is positively correlated to the fundamental exchange rate volatility $\sigma$ but not related to the asset correlation $\rho$ or asset $B$’s volatility $\sigma_{B}$. This observations is consistent with our model implication in Section 4.4.1.
@@ Appendix A Proofs
where the second equality is due to Assumption 3.1-(ii). Because $h(z)$ is strictly decreasing in $z$ and $G(u)$ is strictly decreasing in $u$, we conclude that $h^{\prime}$ is strictly decreasing on $[0,K)$. Moreover, $h^{\prime}(0)=-\beta G(\beta)$ and $\lim_{z\uparrow K}h^{\prime}(z)=-\beta\lim_{u\downarrow 0}G(u)=-\infty$, where the first equality in the above calculation of the limit is the case because it is either the case $K=1$ or the case in which $K<1$ and $\lim_{z\uparrow K}h(z)=-\infty$, and the second equality is due to Assumption 3.1-(ii). Now, denoting the objective function of (89) as $V(d^{A})$, we have
$V^{\prime}(d^{A})=\alpha\beta+(1+f)h^{\prime}(d^{A}).$
Thus, if $G(\beta)\geq\alpha/(1+f)$, $V^{\prime}(d^{A})<0$ for all $d^{A}\in(0,K)$, showing that $V(d^{A})<V(0)=0$ for all $d^{A}\in(0,K)$. If $G(\beta)<\alpha/(1+f)$, there exists a unique maximizer $\hat{d}^{A}\in(0,K)$, which solves the equation $V^{\prime}(\hat{d}^{A})=0$, namely,
@@
On the other hand, dividing both sides of (91) by $\beta$ and recalling Assumption 3.1-(iii) and (90), we conclude that $\hat{d}^{A}$ solves
$F\left(1-\hat{d}^{A},(1-\hat{d}^{A})/G^{-1}\big{(}\alpha/(1+f)\big{)}\right)=F(1,1/\beta),$
which implies that $\hat{d}^{A}$ is strictly increasing in both $\alpha$ and $\beta$ with $G(\beta)<\alpha/(1+f)$ and
$\lim_{\alpha\downarrow G(\beta)(1+f)}\hat{d}^{A}=0,\quad\lim_{\beta\downarrow G^{-1}(\alpha/(1+f))}\hat{d}^{A}=0.$
@@
and the optimal value is strictly larger than 0. Moreover, (i) both $\hat{d}^{A}$ and $\hat{d}^{B}$ are continuous in $(\alpha,\beta)$ with $G(\beta)>\alpha(1+f)$, (ii) $\hat{d}^{A}$ is strictly increasing and $\hat{d}^{B}$ is strictly decreasing in both $\alpha$ and $\beta$ with $G(\beta)>\alpha(1+f)$, and (iii)
$\lim_{\alpha\uparrow G(\beta)/(1+f)}\hat{d}^{i}=0,\quad\lim_{\beta\uparrow G^{-1}(\alpha(1+f))}\hat{d}^{i}=0,\quad i\in\{A,B\}.$
@@
Straightforward calculation yields that
$h(a):=\sup_{c\in(0,1)}H(c,a)=(1+a^{1/\gamma})^{\gamma}$
and the unique maximizer of $H(c,a)$ in $c\in(0,1)$ is $(1+a^{1/\gamma})^{-1}$ when $a>0$. Moreover, $h$ is strictly increasing and $h^{\prime}(a)=\left(\frac{a^{1/\gamma}}{1+a^{1/\gamma}}\right)^{1-\gamma}\leq 1$, $a\geq 0$. As a result,
$\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}=h\mathbb{S}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$
is a contraction mapping on ${\cal X}_{+}$ with Lipschitz constant $(\delta\bar{R}^{1-\gamma})^{N}$. Thus, $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\dots\mathbb{S}_{N-1}$ has a unique fixed point $\tilde{v}_{0}^{*}$ in ${\cal X}_{+}$ and starting from any $\tilde{v}_{0}$, the iterative sequence $(\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\dots\mathbb{S}_{N-1})^{n}\tilde{v}_{0}$, $n=1,\dots,$ converges to $\tilde{v}^{*}$ exponentially. Furthermore, it is straightforward to see that $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$ is monotone, so
@@
In consequence, $\tilde{v}^{*}_{N-1}:=\mathbb{S}_{N-1}\tilde{v}_{0}^{*}\geq\mathbb{S}_{N-1}1>0$, and mathematical induction yields that for every $k=1,\dots,N-2$,
$\tilde{v}^{*}_{k}:=\mathbb{S}_{k}\mathbb{S}_{k+1}\cdots\mathbb{S}_{N-1}\tilde{v}_{0}^{*}=\mathbb{S}_{k}\tilde{v}^{*}_{k+1}>0.$
The standard verification theorem shows that for every $k\in\{0,1,\dots,N-1\}$, $v^{*}_{k}$ as in (56) is the LP’s optimal expected utility for her problem (38), starting from unit wealth at time $k$. Furthermore, the optimal percentage of wealth consumed at time 0 is the maximizer of $H(c,\mathbb{S}_{0}\tilde{v}^{*}_{1})$ in $c\in(0,1)$. Because $\tilde{v}^{*}_{1}>0$ and thus $\mathbb{S}_{0}\tilde{v}^{*}_{1}>0$, the maximizer uniquely exists and is
$c_{0}^{*}=(1+\left(\mathbb{S}_{0}\tilde{v}_{1}^{*}(s,z)\right)^{1/\gamma})^{-1}.$
The optimal percentage of wealth invested in AMM and CEX is
@@
is concave and monotone in $a\geq 0$. As a result,
$\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}=h^{-}\mathbb{S}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$
is a concave, monotone mapping on ${\cal X}_{+}$. It is straightforward to see that
$\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}0=h^{-}(0)=1>0.$
On the other hand, for any constant $L>0$, we have

### v3

@@ 3.3.1 LP’s Investment Return on DEX
$=$ $\frac{p_{k}^{A}y_{(k,1)}^{A}+p_{k}^{B}y_{(k,1)}^{B}}{p_{k}^{A}y_{(k,0)}^{A}+p_{k}^{B}y_{(k,0)}^{B}}\cdot\frac{p_{k}^{A}y_{(k,2)}^{A}+p_{k}^{B}y_{(k,2)}^{B}}{p_{k}^{A}y_{(k,1)}^{A}+p_{k}^{B}y_{(k,1)}^{B}}\cdot\frac{p_{k+1}^{A}y_{(k,2)}^{A}+p_{k+1}^{B}y_{(k,2)}^{B}}{p_{k}^{A}y_{(k,2)}^{A}+p_{k}^{B}y_{(k,2)}^{B}}\cdot\frac{p_{k+1}^{A}y_{(k,3)}^{A}+p_{k+1}^{B}y_{(k,3)}^{B}}{p_{k+1}^{A}y_{(k,2)}^{A}+p_{k+1}^{B}y_{(k,2)}^{B}},$ (23)
where the four terms on the right hand side are denoted as $R^{M}_{(k+1,i)},~{}i=1,2,3,4$, respectively. Here, $R^{M}_{(k+1,1)}$, $R^{M}_{(k+1,2)}$, and $R^{M}_{(k+1,4)}$ stand for the change of the liquidity pool value due to trading by investors and arbitrageurs in Steps 2, 3, and 5, respectively, and $R^{M}_{(k+1,3)}$ stands for the change of the liquidity pool value due to the price shock in Step 4. By (4), (17), (19), and (21), we derive
@@ 3.4.2 Dynamic Programming
    $\delta\bar{R}^{1-\gamma}<1,~{}~{}\text{where }\bar{R}:=\begin{cases}\sup_{(s,z)\in{\cal S}}\sup_{\boldsymbol{\omega}_{k}\in\Omega}U^{-1}\left(\mathbb{E}\Big{[}U\big{(}R^{p}_{k+1}\big{)}\big{]}\mid S_{k}=s,Z_{k}=z\Big{]}\right),~{}\gamma\in(0,1),\\ \inf_{(s,z)\in{\cal S}}\sup_{\boldsymbol{\omega}_{k}\in\Omega}U^{-1}\left(\mathbb{E}\Big{[}U\big{(}R^{p}_{k+1}\big{)}\big{]}\mid S_{k}=s,Z_{k}=z\Big{]}\right),~{}\gamma>1.\end{cases}$ (52)
@@ 3.4.3 The Case of CGMMM Pricing Function
Combining (65)-(68) with (23) and (61)–(64), we derive the return on AMM: $R^{M}_{k+1}/R^{B}_{k+1}$ is a function of $S_{k}$ and random disturbances $(\xi_{k+1},I_{k+1},R^{A}_{k+1}/R^{B}_{k+1})$, but it is independent of $Z_{k}$. Therefore, for the CGMMM pricing function, given the exchange ratio $S_{k}$, the fundamental exchange rate $Z_{k}$ is irrelevant for the LP’s decision. Thus, we can reduce the dimension of the state space.
It is also worth mentioning that for the CGMMM pricing function, the exchange rate ratio takes the form $S_{k}=\frac{\eta}{(1-\eta)}\times\frac{y_{k}^{B}p_{k}^{B}}{y_{k}^{A}p_{k}^{A}},$ where $(y_{k}^{B}p_{k}^{B})/(y_{k}^{A}p_{k}^{A})$ is the relative portfolio weight of asset $B$ to asset $A$ in the liquidity pool. In view of Proposition 3.1, the no-trading region for arbitrageurs can then be written as an interval for this relative portfolio weight. This explains the no-fee interval assumed in Evans et al. (2021).
@@ 4.1 Data Description and Parameter Estimation
We set the length of one period to be 8 hours and assume the LP consumes every 24 hours, so that $N=3$ in our setting. We set the LP’s risk aversion $\gamma=2$. We take ETH and BTC as assets $A$ and $B$, respectively, and estimate the market parameters $\mu_{A}$, $\mu_{B}$, $\sigma_{A}$, $\sigma_{B}$, and $\rho$ using the 8-hour trading data of ETH/USDT and BTC/USDT from July 1st, 2022 to June 30, 2023 obtained from Binance Market Data (https://data.binance.vision). We set $\eta=0.5$ corresponding to the CPMM used in Uniswap, $f=0.5\%$ in the middle of the range of the fee levels on Uniswap v3 (0.05%, 0.3%, and 1%), $\sigma_{I}=0.02$ and $\alpha=0.5$. We summarize the parameter values in Table 1.
@@
Market Parameters LP’s Attributes AMM Parameters
$\delta$ $R_{f}$ $\mu_{A}$ $\mu_{B}$ $\sigma_{A}$ $\sigma_{B}$ $\rho$ $\alpha$ $\sigma_{I}$ $N$ $\gamma$ $\eta$ $f$
0.998 1+0.002% 0.05% 0.038% 1.99% 1.52% 0.8642 0.5 2% 3 2 0.5 0.5%
@@ 4.2.2 Price Slippage
The price slippage is the difference between the marginal exchange rate on AMM and the average exchange rate of trading a positive amount of an asset on AMM. Under CGMMM, it is straightforward to derive that the relative price slippage for acquiring $D_{A}>0$ amount of asset $A$ at marginal exchange rate $E_{0}$ is $\frac{-D^{B}/D^{A}-E_{0}}{E_{0}}=\frac{\frac{1-\eta}{\eta}\left(\big{(}1-D^{A}/y^{A}\big{)}^{\eta/(\eta-1)}-1\right)}{D^{A}/y^{A}},$ where $D^{B}$ is determined by the pricing equation (1), and therefore $-D^{B}/D^{A}$ stands for the average exchange rate. It is straightforward to verify that the above relative price slippage is strictly increasing in the trading size $D^{A}/y^{A}$ and in $\eta$. Similarly, the relative price slippage for acquiring $D_{B}>0$ amount of asset $B$ at marginal exchange rate $E_{0}^{\prime}$ is $\frac{-D^{A}/D^{B}-E_{0}^{\prime}}{E_{0}^{\prime}}=\frac{\frac{\eta}{1-\eta}\left(\big{(}1-D^{B}/y^{B}\big{)}^{(1-\eta)/\eta}-1\right)}{D^{B}/y^{B}},$ which is strictly increasing in the trading size $D^{B}/y^{B}$ and strictly decreasing in $\eta$.
@@ 4.2.3 LP’s Fee Revenue from Liquidity Traders and Loss to Arbitrageurs
$\approx\left[\eta R^{A}(1+f\mathbf{1}_{\tilde{R}^{-1}>(1+f)})\bar{\varphi}^{A}(\tilde{R}^{-1})+(1-\eta)R^{B}(1+f\mathbf{1}_{\tilde{R}^{-1}<1/(1+f)})\bar{\varphi}^{B}(\tilde{R}^{-1})\right]\cdot(y^{A}p^{A}+y^{B}p^{B}),$
@@
Milionis et al. (2024b) consider a continuous-time setting with asset $B$ as the numéraire and thus $R^{B}\equiv 1$. In this setting, $\tilde{R}$ represents the fundamental exchange rate change in an infinitesimally small period of time and thus has a mean of 1, and the expectation of $(\tilde{R}-1)^{2}$ is proportional the fundamental exchange rate volatility. Thus, Milionis et al. (2024b) derive that the LP’s expected loss to arbitrageurs in their setting is proportional to $\eta(1-\eta)$ and the volatility of the fundamental exchange rate; see Example 2 therein.
@@
$\approx\left[\eta\big{(}\left(I^{-1}/(1+f)\right)^{1-\eta}-1\big{)}\mathbf{1}_{I^{-1}>1+f}+(1-\eta)\big{(}\left(I^{-1}(1+f)\right)^{-\eta}-1\big{)}\mathbf{1}_{I^{-1}<1/(1+f)}\right]\cdot(p^{A}y^{A}+p^{B}y^{B})f\xi,$
@@ 4.5 Empirical Test
We use data in 110 AMM pools on Uniswap v3333See https://thegraph.com/hosted-service/subgraph/uniswap/uniswap-v3. and asset price data on Binance [^4] See https://data.binance.vision/. from July 1, 2022 to June 30, 2023. For each AMM pool $j$, we collect its unit trading fee $f_{j}$ from Uniswap v3 and estimate the volatilities of the two assets in the pool $\sigma_{A,j}$ and $\sigma_{B,j}$ and their correlation $\rho_{j}$ using the Binance asset price data. Then we estimate the volatility of the fundamental exchange rate for the pool $j$ as $\sigma_{j}=\sqrt{\sigma_{A,j}^{2}-2\rho_{j}\sigma_{A,j}\sigma_{B,j}+\sigma_{B,j}^{2}}.$ We then run a linear regression between the unit trading fee $f$ and the corresponding fundamental exchange rate volatility $\sigma$, asset correlation $\rho$, and asset $B$’s volatility $\sigma_{B}$ across the 110 AMM pools, and the results are reported in Table 2. At the 5% significance level, the unit trading fee is positively correlated to the fundamental exchange rate volatility $\sigma$ but not related to the asset correlation $\rho$ or asset $B$’s volatility $\sigma_{B}$. This observations is consistent with our model implication in Section 4.4.1.
@@ Appendix A Proofs
where the second equality is due to Assumption 3.1-(ii). Because $h(z)$ is strictly decreasing in $z$ and $G(u)$ is strictly decreasing in $u$, we conclude that $h^{\prime}$ is strictly decreasing on $[0,K)$. Moreover, $h^{\prime}(0)=-\beta G(\beta)$ and $\lim_{z\uparrow K}h^{\prime}(z)=-\beta\lim_{u\downarrow 0}G(u)=-\infty$, where the first equality in the above calculation of the limit is the case because it is either the case $K=1$ or the case in which $K<1$ and $\lim_{z\uparrow K}h(z)=-\infty$, and the second equality is due to Assumption 3.1-(ii). Now, denoting the objective function of (88) as $V(d^{A})$, we have $V^{\prime}(d^{A})=\alpha\beta+(1+f)h^{\prime}(d^{A}).$ Thus, if $G(\beta)\geq\alpha/(1+f)$, $V^{\prime}(d^{A})<0$ for all $d^{A}\in(0,K)$, showing that $V(d^{A})<V(0)=0$ for all $d^{A}\in(0,K)$. If $G(\beta)<\alpha/(1+f)$, there exists a unique maximizer $\hat{d}^{A}\in(0,K)$, which solves the equation $V^{\prime}(\hat{d}^{A})=0$, namely,
@@
On the other hand, dividing both sides of (90) by $\beta$ and recalling Assumption 3.1-(iii) and (89), we conclude that $\hat{d}^{A}$ solves $F\left(1-\hat{d}^{A},(1-\hat{d}^{A})/G^{-1}\big{(}\alpha/(1+f)\big{)}\right)=F(1,1/\beta),$ which implies that $\hat{d}^{A}$ is strictly increasing in both $\alpha$ and $\beta$ with $G(\beta)<\alpha/(1+f)$ and $\lim_{\alpha\downarrow G(\beta)(1+f)}\hat{d}^{A}=0,$ $\lim_{\beta\downarrow G^{-1}(\alpha/(1+f))}\hat{d}^{A}=0.$
@@
and the optimal value is strictly larger than 0. Moreover, (i) both $\hat{d}^{A}$ and $\hat{d}^{B}$ are continuous in $(\alpha,\beta)$ with $G(\beta)>\alpha(1+f)$, (ii) $\hat{d}^{A}$ is strictly increasing and $\hat{d}^{B}$ is strictly decreasing in both $\alpha$ and $\beta$ with $G(\beta)>\alpha(1+f)$, and (iii) $\lim_{\alpha\uparrow G(\beta)/(1+f)}\hat{d}^{i}=0,~{}~{}\lim_{\beta\uparrow G^{-1}(\alpha(1+f))}\hat{d}^{i}=0,~{}~{}i\in\{A,B\}.$
@@
Straightforward calculation yields that $h(a):=\sup_{c\in(0,1)}H(c,a)=(1+a^{1/\gamma})^{\gamma}$ and the unique maximizer of $H(c,a)$ in $c\in(0,1)$ is $(1+a^{1/\gamma})^{-1}$ when $a>0$. Moreover, $h$ is strictly increasing and $h^{\prime}(a)=\left(\frac{a^{1/\gamma}}{1+a^{1/\gamma}}\right)^{1-\gamma}\leq 1$, $a\geq 0$. As a result, $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}=h\mathbb{S}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$ is a contraction mapping on ${\cal X}_{+}$ with Lipschitz constant $(\delta\bar{R}^{1-\gamma})^{N}$. Thus, $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\dots\mathbb{S}_{N-1}$ has a unique fixed point $\tilde{v}_{0}^{*}$ in ${\cal X}_{+}$ and starting from any $\tilde{v}_{0}$, the iterative sequence $(\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\dots\mathbb{S}_{N-1})^{n}\tilde{v}_{0}$, $n=1,\dots,$ converges to $\tilde{v}^{*}$ exponentially. Furthermore, it is straightforward to see that $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$ is monotone, so
@@
In consequence, $\tilde{v}^{*}_{N-1}:=\mathbb{S}_{N-1}\tilde{v}_{0}^{*}\geq\mathbb{S}_{N-1}1>0$, and mathematical induction yields that for every $k=1,\dots,N-2$, $\tilde{v}^{*}_{k}:=\mathbb{S}_{k}\mathbb{S}_{k+1}\cdots\mathbb{S}_{N-1}\tilde{v}_{0}^{*}=\mathbb{S}_{k}\tilde{v}^{*}_{k+1}>0.$ The standard verification theorem shows that for every $k\in\{0,1,\dots,N-1\}$, $v^{*}_{k}$ as in (55) is the LP’s optimal expected utility for her problem (38), starting from unit wealth at time $k$. Furthermore, the optimal percentage of wealth consumed at time 0 is the maximizer of $H(c,\mathbb{S}_{0}\tilde{v}^{*}_{1})$ in $c\in(0,1)$. Because $\tilde{v}^{*}_{1}>0$ and thus $\mathbb{S}_{0}\tilde{v}^{*}_{1}>0$, the maximizer uniquely exists and is $c_{0}^{*}=(1+\left(\mathbb{S}_{0}\tilde{v}_{1}^{*}(s,z)\right)^{1/\gamma})^{-1}.$ The optimal percentage of wealth invested in AMM and CEX is
@@
is concave and monotone in $a\geq 0$. As a result, $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}=h^{-}\mathbb{S}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}$ is a concave, monotone mapping on ${\cal X}_{+}$. It is straightforward to see that $\hat{\mathbb{S}}_{0}\mathbb{S}_{1}\cdots\mathbb{S}_{N-1}0=h^{-}(0)=1>0.$ On the other hand, for any constant $L>0$, we have